A volume of 60.0 mL of a 0.160 M HNO3 solution is titrated with 0.360 M KOH. Calculate the volume of KOH required to reach the equivalence poi
1
Expert's answer
2018-05-02T07:52:08-0400
V(KOH)·c(KOH) = V(HNO3)·c(HNO3)
V(KOH) = V(HNO3)·c(HNO3)/c(KOH) = 60·0.16/0.36 = 26.7 (ml)